OpenStudy (anonymous):

a car with a mass of 2,200 kg can accelerate from 0 to 30 m/s in 6 seconds ? assuming all of the force generated by the engine goes into acceleration of the car ,how much force is the engine producing ?

OpenStudy (anonymous):

That's not a biology question.

OpenStudy (anonymous):

I will tell you how to solve it... calculate acceleration...by using v= u + at v=30m/s, u=0m/s, t=6 sec then multiply the acceleration with the mass of the car.. F= ma... and thats your answer... enjoy